Tohban Report 2014-12-03

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 26 Nov 2014
End Date: 3 Dec 2014
Tohban: Nicole
Tohban email:
Next Tohban:
List all reports


Solar Activity

The sun remained very quiet this week with only 1 M-class flare to report and roughly the same amount of C-class events as last week. The most active region is AR 12222 located just west of disk center in the southern hemisphere, its forecasted to create some further c-class activity. There is a small, not complex, AR on the eastern limb and 7 ARs to the west that will rotate off in the coming days. Next week seems like it will also be quiet.

How many GOES flares occurred?

 Flares above B, C, M, X class were      0    41     1     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      0    24     1     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      0    11     1     0

There were RHESSI flares/GOES flares 305 / 42 over the time range 18-Nov-14 25-Nov-14

Memory Management

SSR emptied at the end of the pass sets for most of the week.


Cold plate temperatures are at 127.6 and 126.2 with their cold tips at 106.6 and 104.9 K, following the normal trend.

Spacecraft Operations

The spacecraft remains in Normal/Vigorous this week and it is recommended to stay in this mode for the coming week. Attenuators, spectrometer, cryocooler and spin are nominal.

Enable fronts for next week.

Data Gaps

A few short VC1 (beginning on 11/28) and VC3 (11/28 and 11/29) data gaps that will be filled in. There were issues with Wallops downlink and transfer of data to BGS MOC, but these should be sorted out by now. There were a few packets that were mislabeled as .bin instead of .dat files.

Detector issues

Detector 7 segmented early this week. We see some decreased counts in the front segment relative to the other detector front segments. David will look at the counts to determine if there is any issue with the reduced front counts. Hazel will check that the thresholds are correct.

Detector 2 front and Detector 9 rear are showing increased noise and reduced livetime due to the light leak.

Other notes

Decimation Normal/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) +/-4 minutes
Night time data (rears) full nighttime events
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Nominal
Detector problems? Detector 2 front and D9 rear light leak.
Personal tools